Suppose your expectations regarding the stock market are as follows:

State of the Economy Probability HPR
Boom 0.3 42%
Normal growth 0.4 15
Recession 0.3 -18

What is the mean?

What is the Standard Deviation?
Answer #1

Mean=Respective HPR*Respective probability

=(0.3*42)+(0.4*15)+(0.3*-18)=13.2%

probability HPR probability*(HPR-Mean)^2
0.3 42 0.3*(42-13.2)^2=248.832
0.4 15 0.4*(15-13.2)^2=1.296
0.3 -18 0.3*(-18-13.2)^2=292.032
Total=542.16%

Standard deviation=[Total probability*(HPR-Mean)^2/Total probability]^(1/2)

=[542.16]^(1/2)

=23.28%(Approx).